John is about as old in days as Jake is in weeks.
John is also as many months as Jim is in years.
Together, the 3 of them are 140 years old.
How old is Jim?

Answer by ankor@dixie-net.com(22740)

John is about as old in days as Jake is in weeks.
John is also as many months as Jim is in years.
Together, the 3 of them are 140 years old.
How old is Jim?
:
Let the three guys be x, y, z respectively
x = john's age (all in yrs)
y = Jake's age
z = Jim's age
:
John is about as old in days as Jake is in weeks
365x = 52y
Divide both sides by 52
7x = y
:
John is also as many months as Jim is in years.
12x = z
:
Together, the 3 of them are 140 years old.
x + y + z = 140
replace y with 7x, replace z with 12x
x + 7x + 12x = 140
20x = 140
x = 140/20
x = 7 yrs old is John
:
Find Jim
z = 12(7)
z = 84 yrs old is Jim
:
Check this by finding the sum of the 3 ages